In a group of 15, 7 have studied Latin, 8 have studied Greek, and 3 have not studied either. How many of these studied both Latin and Greek?

a) 0

b) 3

c) 4

d) 5
Answer #1

Let A = Latin and B = Greek

n(A or B) = 15 - n(A or B)'

= 15 - 3

= 12

n(A or B) = n(A) + n(B) - n(A and B)

12 = 7 + 8 - n(A and B)

n(A and B) = 3